OpenStudy (anonymous):

Help please! A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (anonymous):

C = 0.765x + 0.06(0.765x) x = 100 plug in

OpenStudy (anonymous):

so its c=0.765(100)+0.06(0.765x100) ??

OpenStudy (anonymous):

yes, what do u get?

OpenStudy (anonymous):

c= 76.5+4.59

OpenStudy (anonymous):

correct, one more step

OpenStudy (anonymous):

is it 81.09?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much!(:

OpenStudy (anonymous):

you welcome